### v1

# FocusLLM: Scaling LLM’s Context by Parallel Decoding

Abstract
Empowering LLMs with the ability to utilize useful information from a long context is crucial for many downstream applications. However, achieving long context lengths with the conventional transformer architecture requires substantial training and inference resources. In this paper, we present FocusLLM, a framework designed to extend the context length of any decoder-only LLM, enabling the model to focus on relevant information from very long sequences. FocusLLM processes long text inputs by dividing them into chunks based on the model’s original context length to alleviate the issue of attention distraction. Then, it appends the local context to each chunk as a prompt to extract essential information from each chunk based on a novel parallel decoding mechanism, and ultimately integrates the extracted information into the local context. FocusLLM stands out for great training efficiency and versatility: trained with an 8K input length with much less training cost than previous methods, FocusLLM exhibits superior performance across downstream long-context tasks and maintains strong language modeling ability when handling extensive long texts, even up to 400K tokens. Our code is available at https://github.com/leezythu/FocusLLM.
FocusLLM: Scaling LLM’s Context by Parallel Decoding
Zhenyu Li1, Yike Zhang1, Tengyu Pan1, Yutao Sun1, Zhichao Duan1, Junjie Fang2, Rong Han1, Zixuan Wang1, Jianyong Wang1thanks: Corresponding author 1Tsinghua University 2Xiamen University
1 Introduction
The importance of extending the context length of large language models (LLMs) cannot be overstated. In numerous applications, ranging from complex document analysis to generating coherent long-form text, the ability to effectively utilize extended context is critical. For instance, in tasks such as document summarization and question answering over lengthy articles, a more extensive context allows for a more comprehensive understanding and accurate responses Li et al. (2024a). However, leveraging long contexts in LLMs presents several formidable challenges. (1) The computational complexity of transformers Vaswani et al. (2017) grows quadratically with the sequence length, rendering the training process prohibitively expensive. (2) LLMs exhibit poor extrapolation performance for longer sequences, even after additional fine-tuning Chen et al. (2023a); Peng et al. (2023). (3) Acquiring high-quality long-text datasets, which are essential for training and fine-tuning, is exceedingly difficult Xiong et al. (2023); Wang et al. (2022).
Figure 1: A comparison between FocusLLM and previous context scaling methods on the passkey retrieval task, including CEPE Yen et al. (2024), LongLLaMA Tworkowski et al. (2024) and Activation Beacon Zhang et al. (2024a). Our method extrapolates beyond the original context length of LLaMA, achieving 99% accuracy at a context length of 400K, with less training cost.
To circumvent the substantial costs of directly scaling the window length by fine-tuning on longer inputs, many approaches have attempted to modify the attention mechanism Xiao et al. (2023); Han et al. (2023) or compress tokens Zhang et al. (2024a); Chevalier et al. (2023); Ge et al. (2023) to theoretically achieve infinite length. While these methods can maintain lower perplexity over extended texts, the loss of information from earlier parts of the text hampers the model’s ability to perform precise understanding tasks such as information verification or question answering Zhang et al. (2024a).
Figure 2: The standard autoregressive model architecture. Suppose the input consists of a lengthy segment of information (e.g., a long document) and a user instruction (e.g., a question), with the total length being $L$. Due to the attention mechanism, the computational complexity is $O(L^{2})$, and $L$ is constrained by the length to which the model has been pre-trained. The performance on texts exceeding the length $L$ tends to deteriorate sharply.
Figure 3: The FocusLLM framework. A small fragment of the local context is appended to each chunk (where $k$ = 3), which consists of memory tokens. The representations of the candidate tokens, obtained through parallel decoding, are then concatenated and integrated as memory back into the local context.
In this paper, we propose a training efficient and effective solution that enables large language models (LLMs) to focus on relevant content within long-context documents and generate high-quality outputs, called FocusLLM, by utilizing parallel decoding. FocusLLM is designed to adapt to any decoder-only language model with the modifications: i) the original model parameters are frozen to retain their generalization capabilities; ii) A small number of trainable parameters are added to enable the model to aggregate information from parallel decoding of different chunks. Not only does FocusLLM maintain low perplexity on long texts, but it also performs well on downstream tasks requiring precise understanding. Our evaluation on a widely adopted passkey retrieval task (Figure 1) has shown that FocusLLM, in contrast to the recently proposed context scaling methods, can sustain the accuracy of text comprehension on much longer texts.
FocusLLM bears several standout features: (1) Length Scaling: FocusLLM breaks the inherent positional limitations, allowing the model to handle text lengths that are expanded by tens or even hundreds of times. (2) Training Efficiency: Unlike full fine-tuning, we keep the original model parameters frozen and add only a small number of trainable parameters. The training process can be completed within a 0.5B tokens training budget, which is significantly smaller than that of previous works. (3) Versatility: FocusLLM not only excels in downstream tasks with clear instructions, such as question answering, but also demonstrates strong language modeling capabilities over long documents.
We employ the FocusLLM framework to LLaMA-2-7B Touvron et al. (2023b), which has a default context length of 4K. To validate the effectiveness of FocusLLM, we evaluate it across a variety of tasks. Initially, we assessed FocusLLM’s language modeling capability. Trained with only 8K input length, FocusLLM maintains low perplexity on documents comprising 128K tokens and even longer sequences. Subsequently, to comprehensively evaluate the applicability of FocusLLM in real-world scenarios, we utilized two widely used benchmarks: Longbench Bai et al. (2023) and $\infty$-Bench Zhang et al. (2024b). Longbench encompasses a diverse range of tasks, while $\infty$-Bench requires models to perform inference on extremely long sequences (>100K tokens). Experimental results demonstrate that FocusLLM has achieved superior performance on both benchmarks, surpassing all baselines including length extrapolation models, continual training models, and similar models designed for extreme long sequences.
In conclusion, FocusLLM is a training-efficient framework capable of achieving effective understanding and reasoning on long sequences at a minimal cost. We hope that FocusLLM can contribute to research related to developing easy-to-implement long-context capabilities.
2 Methodology
In this section, we introduce the design methodology of FocusLLM. First, we describe how we construct FocusLLM based on the architecture of LLM to enable it to handle extremely long text contexts. Then, we explain the training process of FocusLLM.
2.1 Architecture
As illustrated in Figure 2, the standard model architecture has a quadratic complexity and a corresponding limited context length. This limitation restricts the model’s application to longer texts, and FocusLLM is designed to address these drawbacks.
The overall framework of FocusLLM is simple and intuitive. Each decoder in Figure 3 shares the same model (e.g. LLaMA-2). Besides, for the decoder handling each chunk, we augment the original decoder with a small set of additional parameters.
2.1.1 Notations.
Given a long sequence with $S$ tokens $\{x_{1},...,x_{S}\}$, we segment them into memory tokens $\{x_{1},...,x_{m}\}$ and local tokens $\{x_{m+1},...,x_{S}\}$, with the length of local tokens not exceeding the model’s default context length, denoted as $L$. Concurrently, we divide the memory into chunks, labeled as $C_{1},C_{2},...,C_{k}$, with each chunk’s size also not exceeding $L$. These chunks can represent distinct documents or constitute a single long document.
We define the original decoder model as $F_{dec}$ and its hidden dimension $d_{dec}$. When processing the memory, to endow the model with the capability to generate candidate tokens, we introduce a small set of new parameters, resulting in the modified model $F^{\prime}_{dec}$. The candidate token is denoted as the trainable hidden states corresponding to the last local token $x_{S}$ in each chunk, serving as a signal indicating whether this chunk contains information relevant to the local context, as well as facilitating the prediction of the next token based on the current chunk.
2.1.2 Local Context Injection.
The aim of our FocusLLM is to distribute the burden of understanding long texts across each chunk. Therefore, unlike previous approaches that encode memory tokens and save them as cache, we append a small fragment of local tokens (we refer to it as the ‘prompt’ in the figure) behind each chunk and perform parallel decoding within each chunk, as shown in Figure 3. This strategy allows for a more efficient handling of long sequences by focusing the computational effort on relevant segments of the text, while maintaining the global information necessary for accurate decoding. We can formally define this process as follows:
$\hat{C_{i}}\leftarrow\{C_{i};x_{m+j},...,x_{S}\}\ \ i=1,...,k;1\leq j\leq S-m$ (1)
Here $j$ is a hyperparameter that determines the number of local tokens appended to each chunk. To reduce computational overhead, there is no need to append all local tokens. For instance, if the local context length is 2K, we could concatenate only 512 tokens or just a short instruction. In experiments, we adopt a default length of 512 tokens for inference, which is sufficient to encapsulate the necessary local contextual information.
2.1.3 Parallel Decoding.
FocusLLM transforms the next token prediction process based on long sequences into the simultaneous generation of candidate tokens from different chunks, followed by aggregating these candidate tokens as memory to produce the final token. We refer to this process as parallel decoding.
Differences from previous encoding methods. Here, we will elucidate the meaning of the term ‘decoding’ and explain the logical differences from previous encoding methods. Many previous methods involve encoding the context into some form of memory and retaining the key-value (KV) cache for subsequent use Yen et al. (2024); Tworkowski et al. (2024); Zhang et al. (2024a). For instance, CEPE Yen et al. (2024) divides long texts into chunks and then encodes each chunk with a smaller language model. FocusLLM also divides long texts into chunks, but within each chunk, FocusLLM undergoes a decoding process.
Specifically, the purpose of encoding is to transform a segment of text into hidden states for later use, whereas decoding is about generating the next token based on the current text. From this perspective, our method can be viewed as a form of implicit decoding. This is because our objective is not to acquire representations for entire chunks of text but rather to maintain information about the next token (based on the current chunk $\hat{C_{i}}$) via a special candidate token. Although we do not explicitly map the hidden states of the special candidate tokens through a softmax layer to the corresponding tokens in the vocabulary, our approach is fundamentally aligned with the concept of decoding. For further discussion, please refer to Appendix A.
How to obtain the representations of candidate tokens. Inspired by Zhang et al. (2024a), in order to preserve the generalizability of the original model as much as possible, we only add a new set of trainable parameters to the linear projection matrices of each layer:
$\{W^{\prime}_{Q},W^{\prime}_{K},W^{\prime}_{V},W^{\prime}_{O}\}_{l}$ (2)
where $W^{\prime}_{Q}$, $W^{\prime}_{K}$, $W^{\prime}_{V}$, and $W^{\prime}_{O}$ represent the new parameters for the query, key, value, and output matrices, respectively, and $l$ denotes the layer number. The modified decoder model are denoted as $F^{\prime}_{dec}$.
We apply parallel decoding on each chunk and obtain one candidate token from every chunk:
$e_{i}=F^{\prime}_{dec}(\hat{C_{i}})$ (3)
where $e_{i}$ consists of key-value hidden states $K_{e}$ and $V_{e}$ of the last token in each layer. More formally, the output of the candidate token in the self-attention module is calculated similarly to Zhang et al. (2024a):
$Q_{e}\leftarrow HW^{\prime}_{Q}\quad K_{e}\leftarrow HW^{\prime}_{K}\quad V_{e}\leftarrow HW^{\prime}_{V}\quad$ (4)
$A\leftarrow softmax\left(Q_{e}\left(K\oplus K_{e}\right)^{T}\right)$ (5)
$O_{e}\leftarrow V_{e}{W^{\prime}_{O}}^{T}\quad V_{e}\leftarrow A\left(V\oplus V_{e}^{T}\right)$ (6)
where $H\in R^{d_{dec}}$ is the input hidden states of the candidate token, $\oplus$ represents the concatenation of matrices, and $K,V$ correspond to the representations of the normal tokens in one chunk. Finally, the generated candidate tokens are concatenated with the local tokens and are subsequently processed by a frozen decoder. To optimize the candidate tokens, we design two loss functions of language modeling in Section 2.3. Note that the process of obtaining the candidate token from each chunk is independent, which allows us to perform parallel forwarding for each chunk. This is the reason we refer to this process as ‘parallel decoding’.
2.1.4 Efficiency.
The parallel decoding mechanism of FocusLLM effectively reduces the computational complexity of the standard architecture. Specifically, when dealing with very long sequences, the primary computational burden in the transformer architecture lies in the attention mechanism, which has a complexity of $O(L^{2})$, where $L$ represents the total sequence length. By dividing the sequence into $n$ chunks, the complexity within each chunk becomes $O((L/n)^{2})$. Therefore, when we process chunks in parallel, the time complexity can be reduced to $O((L/n)^{2})$. And the space complexity of $n$ chunks becomes approximately $O((L/n)^{2}*n)=O(L^{2}/n)$. This means that compared to a standard transformer, FocusLLM can reduce the computational complexity to a fraction, $1/n$ or even more of the original theoretically, where $n$ is the number of chunks into which the sequence is divided. In experiments, the longer the sequence length, the more apparent the improvement in efficiency.
2.2 Data
To ensure the generalizability of our method and to maintain fairness in comparison with the baselines, we leverage the same training corpus as Activation Beacon Zhang et al. (2024a), with sequence lengths varying between 3K and 8K tokens. Specifically, we utilized 70K samples from the pre-training dataset RedPajama Together (2023b) and 10K samples from the instruction-following dataset LongAlpaca Chen et al. (2023b). RedPajama is an open-source pre-training dataset for LLaMA-1 Touvron et al. (2023a), while LongAlpaca is an instruction-following dataset for question answering. Both datasets are widely utilized in previous work. The detailed statistics of our training data are reported in Table 1.
Table 1: Length distribution of training corpus.
Length 2Ksimilar-$\sim$4K 4Ksimilar-$\sim$6K 6Ksimilar-$\sim$8K Total
Count 30K 16K 34K 80K
Portion 38% 20% 42% 100%
2.3 Training
We primarily conduct experiments on the LLaMA2-7B-Chat model. After incorporating the parameters mentioned in Section 2.1, the additional parameters amount to only 2B approximately. We train the model using an auto-regressive approach:
Auto-Regressive Loss. FocusLLM is trained using a natural auto-regressive method. Specifically, we train the model to predict the next token, which encourages the candidate token to aggregate useful information from each chunk. Therefore, the loss is only applied to the local tokens:
$\min_{F^{\prime}_{dec}}-\sum_{i=2}^{L}\log(p(x_{i}\mid e_{1},\ldots,e_{k},x_{1},\ldots,x_{i-1};F^{\prime}_{dec}))$ (7)
In Section 2.1.1, we noted that local tokens are positioned subsequent to the memory tokens. However, to explicitly train the model to harness its ability to recover information from the memory tokens, we also set the local tokens to be a substring of the memory tokens. Specifically, based on the different selection methods for local tokens, we design two types of loss functions for joint training. i) Firstly, if the last $L$ tokens from a long document are selected as local tokens, with the remainder serving as memory tokens, we term this loss the ‘Continuation loss’, as it trains the model’s ability to naturally continue generating new tokens based on the context. ii) Alternatively, if we take the entire long document as memory and then randomly select $L$ continuous tokens from it as local tokens, we define this loss the ‘Repetition loss’, because it trains the model’s ability to repeat when clear information from the context is already available. Subsequent experiments have proven that both types of loss are important.
Generalizing Chunk Size. To ensure the model exhibits robust generalizability across various chunk sizes and number of candidate tokens, we maintain a constant local context size of 2048, while the chunk size is randomly selected from the set {64, 128, 256, 1024, 2048} during training.
Table 2: Language Modeling Assessment: perplexity analysis of various context scaling methods on the PG19, Proof-Pile, and CodeParrot. FocusLLM successfully extends context of the original Llama model and maintains low perplexity on extremely long sequences.
PG19 Proof-Pile CodeParrot
Method 4K 16K 32K 100K 4K 16K 32K 100K 4K 16K 32K 100K
Llama-2-7B 9.21 ${>}10^{3}$ ${>}10^{3}$ OOM 3.47 ${>}10^{3}$ ${>}10^{3}$ OOM 2.55 ${>}10^{3}$ ${>}10^{3}$ OOM
PI 9.21 19.5 ${>}10^{2}$ OOM 3.47 5.94 33.7 OOM 2.55 4.57 29.33 OOM
NTK 9.21 11.5 37.8 OOM 3.47 3.65 7.67 OOM 2.55 2.86 7.68 OOM
StreamingLLM 9.21 9.25 9.24 9.32 3.47 3.51 3.50 3.55 2.55 2.60 2.54 2.56
AutoCompre.-6K 11.8 ${>}10^{2}$ ${>}10^{3}$ OOM 4.55 ${>}10^{2}$ ${>}10^{3}$ OOM 5.43 ${>}10^{2}$ ${>}10^{3}$ OOM
YaRN-128K 6.68 6.44 6.38 OOM 2.70 2.47 2.41 OOM 2.17 2.04 2.00 OOM
LongChat-32K 9.47 8.85 8.81 OOM 3.07 2.70 2.65 OOM 2.36 2.16 2.13 OOM
LongAlpaca-16K 9.96 9.83 ${>}10^{2}$ OOM 3.82 3.37 ${>}10^{3}$ OOM 2.81 2.54 ${>}10^{3}$ OOM
LongLlama 9.06 8.83 OOM OOM 2.61 2.41 OOM OOM 1.95 1.90 OOM OOM
Activation Beacon 9.21 8.54 8.56 8.68 3.47 3.42 3.39 3.35 2.55 2.54 2.53 2.55
FocusLLM 9.21 9.19 9.17 10.59 3.47 3.17 3.43 2.57 2.55 2.01 2.27 3.02
3 Experiments
In this section, we will conduct a comprehensive evaluation of the effectiveness of FocusLLM, spanning both language modeling and a variety of downstream tasks.
3.1 Experimental Details
We aligned most of our experimental settings with those of Activation Beacon Zhang et al. (2024a) to ensure comparable results. Specifically, we conducted training on a Linux server equipped with 8×A100 GPUs, each with 40GB of memory. The training was carried out for 10,000 steps, equivalent to one epoch of the entire training dataset, using a batch size of 8 and a learning rate of 5e-5 with a linear scheduler. To conserve GPU memory, we employed deepspeed’s zero2_offload optimizing stage. The training process was completed in approximately 20 hours.
For hyper-parameters, during training, the chunk size was randomly selected from the set {64, 128, 256, 1024, 2048}. For the length of tokens injected into each chunk, we set a default of 512 tokens for inference. And we ensured this length did not exceed the chunk size in the training procedure. As a result, the length of injected tokens was $\min\{512,chunk\ size\}$.
3.2 Long-context Language Modeling
In this section, we evaluate the fundamental capabilities of FocusLLM on long-context language modeling benchmarks, with text lengths ranging from 4K to 128K tokens.
Datasets. We perform the evaluation on three datasets: PG19 Rae et al. (2019), Proof-Pile Azerbayev et al. (2023), and CodeParrot Tunstall et al. (2022). These three datasets encompass 100 long test cases related to books, arXiv papers, and code repositories, respectively. The results of baseline models are token from Zhang et al. (2024a) for comparison. Following the setting of Yen et al. (2024), as FocusLLM relies on the last decoder to perform generation, we calculate the perplexity on the last 256 tokens of each sequence, and for the 128K length, we filter out documents exceeding 128K tokens and evaluate 10 samples due to data scarcity and computational cost.
Model. FocusLLM is based on LLaMA-2-7B (chat), hence the models for comparison are all on the same scale, 7B. The baseline models can be categorized into the following types: i) Methods focusing on the modification of positional encoding, including Positional Interpolation Chen et al. (2023a), the NTK-Aware Scale ROPE [^1] https://www.reddit.com/r/LocalLLaMA/comments/14lz7j5/
ntkaware_scaled_rope_allows_llama_models_to_have/
, and the training-free method StreamingLLM Xiao et al. (2023), which is based on attention sinks. ii) Fine-tuned methods trained on long inputs, such as LongAlpaca-16K Chen et al. (2023b), LongChat-32K Li et al. (2023a), and YaRN-128K Peng et al. (2023). iii) Methods with designed structures specifically for long contexts, including AutoCompressor-6K Chevalier et al. (2023), LongLlama Tworkowski et al. (2024) and Activation Beacon Zhang et al. (2024a). For instance, Activation Beacon achieves compression of long texts by training the model to represent the information of a regular text segment with a small number of beacon tokens.
Analysis. The results are presented in Table 2. Here are several observations we can make: (1) Compared to the basic LLaMA-2-7B model and some fine-tuning free methods, our model demonstrates superior performance. When extending the context length from 4K to longer, the perplexity becomes lower, indicating that information from a longer context can be effectively utilized. (2) There is a slight increase in the perplexity at longer lengths on CodeParrot. One possible reason is that as the number of candidate tokens from parallel decoding in the memory increases, the model’s output probabilities become more evenly distributed. However, this does not affect the effectiveness in downstream tasks (as discussed in Section 3.3). (3) FocusLLM achieves comparable performance to fine-tuned full-attention methods. This result is notable because our model operates with significantly higher training efficiency. For instance, LongLlama is fine-tuned using 7B tokens with all parameters being trainable. In contrast, FocusLLM uses $1/10$ of the training budget and $1/3$ of the parameters. (4) FocusLLM can maintain language modeling capabilities at lengths much longer than other models while retaining precise comprehension of the entire text. In experiments, FocusLLM can maintain stable perplexity at lengths of 400K, which means a 100-fold expansion of the original LLaMA-2-7B’s context length. Although models like StreamingLLM and Activation Beacon can still achieve lower perplexity by compressing tokens, they are unable to recover the previous context information, which severely affects their capabilities in downstream tasks.
In summary, FocusLLM achieves comparable language modeling performance on shorter sequences with a small training cost, while also being capable of handling much longer texts. Building on the design of parallel decoding, FocusLLM avoids the information loss associated with the preceding context, which is a common issue in compression models. We will provide a detailed demonstration of the downstream task performance in the subsequent sections.
Table 3: The results of Vicuna-based models on $\infty$-Bench and LongBench. The models on the right part can process extremely long inputs. On both benchmarks, FocusLLM achieves significant improvements compared to strong baselines.
Vicuna-7B-v1.5 (4K)
Original LChat Vic-16K Yarn-128K PI NTK Stream InfLLM FocusLLM
$\infty$-Bench (214k tokens) Math.Find 11.71 9.43 13.43 17.14 OOM OOM 6.00 11.14 11.71
En.MC 30.13 24.45 34.06 27.95 OOM OOM 32.31 31.44 32.31
Code.Debug 38.83 27.66 35.03 22.59 OOM OOM 46.19 34.26 28.43
Retrieve.KV 1.40 1.40 1.00 0.00 OOM OOM 0.00 0.60 12.40
Retrieve.Number 4.41 23.90 10.34 56.61 OOM OOM 4.41 81.69 83.56
Retrieve.PassKey 5.08 28.64 15.25 92.71 OOM OOM 4.92 99.15 95.76
Average 15.26 19.25 18.19 36.17 15.64 43.05 44.03
LongBench (31K tokens) NarrativeQA 11.19 20.35 17.85 19.67 0.78 5.66 15.61 15.53 21.14
Qasper 13.79 29.35 25.85 11.10 2.71 21.17 23.84 23.57 31.07
MultiFieldQA 22.08 42.55 37.15 35.06 1.01 36.76 32.80 37.14 36.73
HotpotQA 12.71 33.19 24.72 11.94 1.35 19.54 22.17 22.53 40.65
2WikiMQA 13.99 24.33 21.41 12.02 1.17 14.51 18.38 18.82 20.30
Musique 4.81 14.71 8.44 7.52 0.71 4.30 6.30 5.24 14.20
GovReport 27.67 30.83 27.62 29.46 1.9 25.26 23.18 26.79 26.66
QMSum 19.72 22.93 22.63 21.53 1.29 19.48 20.09 20.91 20.50
MultiNews 26.61 26.63 27.88 16.04 1.16 25.88 26.19 26.43 27.45
TREC 69.00 66.50 69.00 68.50 4.50 59.00 61.00 67.50 68.00
TriviaQA 81.94 83.99 85.63 88.21 0.90 25.85 78.81 84.36 81.63
SAMSum 35.12 12.83 9.15 26.52 0.12 5.05 32.46 31.89 35.36
PassageRetrieval 9.00 30.50 4.00 16.25 0.62 5.00 6.00 9.00 15.67
LCC 64.53 54.79 50.64 66.39 21.54 53.65 63.70 61.41 62.79
RepoBench-P 50.17 58.99 44.94 55.82 19.36 44.58 48.26 47.52 53.72
Average 30.82 34.70 31.79 32.40 3.94 24.38 31.92 33.24 36.17
3.3 Downstream Tasks
Datasets. To assess the capabilities of FocusLLM in real-world scenarios, we select two widely used datasets: Longbench Bai et al. (2023) and $\infty$-Bench Zhang et al. (2024b). Longbench offers an evaluation on a variety of tasks including question answering, summarization, few-shot learning, mathematical counting, and code completion. The average and $95$% percentile length are 12.8K and 31K tokens, respectively. $\infty$-Bench is designed to test a model’s ability to understand and reason over super long contexts. The average length and the $95$% percentile length are 145.1K and 214K tokens. For more detailed dataset statistics, please refer to Appendix B. We believe that these two benchmarks can comprehensively reflect the capabilities of the model on downstream tasks. For evaluations on the Longbench, we adopt a larger local context size of 3,500 tokens for FocusLLM, consistent with the official setting.
Models. The purpose of our paper is to enable LLMs with limited context length to understand extremely long sequences at a very minimal cost. Therefore, in addition to comparing FocusLLM with the three types of baselines mentioned in Section 3.2, we mainly focused on comparing FocusLLM with recent proposed models capable of processing extremely long streaming inputs. Specifically, StreamingLLM utilizes a sliding window mechanism; InfLLM Xiao et al. (2024) stores processed context into memory units and retrieves it using attention scores; Activation Beacon compresses the preceding text to maintain a smaller context length. CEPE Yen et al. (2024) adopts a small encoder to process long inputs chunk by chunk and feeds the memory to a decoder by cross-attention.
Main Results. The experimental results are displayed in Table 3 and 4. We reference some baseline results from Xiao et al. (2024), which are based on the Vicuna-7B-v1.5 model. Vicuna-7B-v1.5 is based on LLaMA-2-7B but fine-tuned on conversational data. For a fair comparison, we also train a Vicuna version of FocusLLM. For YaRN-128K, we select the version based on Mistral-7B-inst-v0.2, which is stronger than Vicuna. For LongLlama, as they do not have a version based on the Llama2, we directly utilize the officially released model. Details regarding the effective lengths for each method presented in Tables 3 and 4 are elaborated upon in Appendix C.
From the experimental results, we can make the following observations: (1) FocusLLM outperforms all baseline models, achieving the best results on both the relatively shorter benchmark Longbench and the extremely long benchmark $\infty$-Bench. This demonstrates FocusLLM’s capability for effective understanding and reasoning on long sequences and its broad applicability. (2) Different types of baseline models exhibit various shortcomings. For training-free models like PI and NTK, extending the length to 128K comes with a significant sacrifice in performance. Due to the lack of precise understanding of the full context, models that employ sliding window or condensing techniques, such as StreamingLLM and Activation Beacon perform poorly on $\infty$-Bench, with performance nearly approaching zero on some tasks. As for fine-tuned models like LongChat and CEPE, their limitation is the restricted supported length. For example, CEPE struggles to handle lengths beyond 128K effectively Yen et al. (2024). (3) The approaches of length extrapolation and continual training on long inputs, while capable of scaling context length, introduce substantial computational and memory costs. Consequently, this can lead to exceeding GPU memory limits when dealing with long sequences. In contrast, FocusLLM processes the text in chunks and utilizes parallel decoding, which significantly conserves both the memory and time for inference.
In summary, compared to previous context scaling methods, FocusLLM achieves superior performance across a variety of downstream tasks at the same parameter scale, and can be deployed with limited resources.
Table 4: The results of LLaMA2-based models on tasks of LongBench. L_L represents Long Llama and A_B represents Activation Beacon. FocusLLM outperforms recently proposed memory-based and compression-based methods, and maintains attention to all tokens of context.
Llama-7B-chat (4K)
Original CEPE L_L A_B FocusLLM
NarrativeQA 18.70 22.14 - - 20.38
Qasper 19.20 26.34 - - 21.73
MultiFieldQA 36.80 31.56 - - 36.91
   -Average 24.90 26.68 30.12 27.14 26.34
HotpotQA 25.40 34.95 - - 38.95
2WikiMQA 32.80 32.39 - - 32.95
Musique 9.40 9.76 - - 15.39
   -Average 22.60 25.70 16.37 28.28 29.10
GovReport 27.30 13.90 - - 25.54
QMSum 20.80 20.30 - - 21.86
MultiNews 25.80 3.10 - - 26.35
   -Average 24.70 12.43 24.19 25.15 24.55
TREC 61.50 68.50 - - 68.00
TriviaQA 77.80 87.90 - - 85.08
SAMSum 40.70 32.38 - - 41.63
   -Average 60.00 62.92 60.31 60.72 64.81
LCC 52.40 66.21 - - 58.42
RepoBench-P 43.80 58.94 - - 54.27
   -Average 48.10 62.57 66.05 57.83 56.35
Average 35.20 36.31 37.50 38.54 39.01
4 Further Exploration
In this section, we conduct further explorations of FocusLLM. Initially, we investigate the upper limits of the sequence lengths that FocusLLM can effectively handle. Subsequently, we perform a quantitative analysis of several key parameters within the framework. Ultimately, we undertake ablation studies on the training loss functions to discern their impacts on the model’s overall performance.
4.1 Scaling to 400K Context
We contend that FocusLLM is capable of processing extremely long sequences. To validate this, we first conduct experiments on the passkey retrieval task Mohtashami and Jaggi (2024). The results, as illustrated in Figure 1, demonstrate that FocusLLM maintains nearly 100% effectiveness at lengths of up to 400K, outperforming all other models. In addition to this, we extended the language modeling experiments introduced in Section 3.2 to 400K, a length at which most models fail to manage effectively. The outcomes of these extended experiments are presented in Figure 6.
To conclude, FocusLLM’s exceptional performance in both passkey retrieval and language modeling tasks, even at the substantial length of 400K tokens [^2] Constrained by hardware, the maximum length we are able to test is 400k tokens., indicates that it can achieve good performance in the vast majority of real-world scenarios, which underscores the effectiveness and versatility of FocusLLM.
Figure 4: Perplextiy on the PG19 dataset of FocusLLM compared to methods PI and NTK. FocusLLM can maintain low perplexity even at a length up to 400K tokens.
Figure 5: Memory usage of FocusLLM compared against previous long context methods. FocusLLM without parallel forwarding exhibits a linear growth pattern.
Figure 6: Comparison of inference time. The time taken by FocusLLM is superior to other long-context methodologies but slightly higher than the standard approach.
4.2 Memory Footprint and Inference Time
For models that focus on long-form text, aside from training costs, another critical aspect is the memory footprint and inference time, especially as sequence lengths increase. In this section, we compare FocusLLM with several previous long-context methods capable of retaining global information by preserving the cache of all context: Standard (PI/NTK), LongLlama, and CEPE. The results are shown in Figure 6 and Figure 6. As for models like Activation Beacon and StreamingLLM, although they maintain a constant memory footprint by only retaining cache for a fixed window, they struggle with the precise understanding of extremely long texts. Therefore, they are not the primary subjects of comparison.
In the Figure, "FocusLLM with or without parallel" signifies that we process each chunk either concurrently or sequentially. The findings indicate that: (1) Serial processing is good enough in terms of memory usage and inference time. When ample memory resources are available, parallel processing is more efficient. And in practice, the parallel level is also controllable, making our model robust and applicable under most scenarios.(2) Although FocusLLM splits long texts into numerous chunks, resulting in a slightly longer inference time compared to the standard approach, it holds a significant advantage over other long-context methods.
4.3 Chunk Size
We conduct an investigation into the impact of different chunk sizes on performance. In theory, larger chunk sizes, as long as they do not exceed the model’s default context length (e.g., 4K for LLaMA-2), are preferable because they allow for processing the memory with a smaller number of forward passes. However, smaller chunk sizes may enable more precise processing.
In experiments, we maintain a total sequence length of 8K, testing the perplexity using different chunk sizes on the same samples of PG19. We select {256, 512, 1024, 2048} as our test sizes. The results are shown in Figure 7. We observe that there is no consistent trend in perplexity as the chunk size increases; it remains relatively stable. This confirms our hypothesis that we can employ larger chunk sizes on models with longer default context lengths (e.g. LLaMA-2-32K). We will explore this direction in our future work.
Table 5: Investigations into the training loss function and local context size of FocusLLM (based on LLaMA2-7B). We present the results for representative tasks from LongBench and $\infty$-Bench. For instance, NarrativeQA belongs to Single-Doc QA, while TREC relates to Few-shot learning. The Hyper Params is denoted as (local context size, chunk size).
LongBench $\infty$-Bench
Hyper Params. NarrativeQA TREC Math.Find En.MC Retrieve.PassKey
FocusLLM (2K, 2K) 18.53 65.5 13.43 31.00 99.32
Continuation Loss only (2K, 2K) 17.36 60.5 13.71 27.95 1.69
Repetition Loss only (2K, 2K) 17.05 62.0 12.86 26.64 91.19
Local Context Size $\downarrow$ (1K, 2K) 17.87 63.0 8.86 29.69 99.32
4.4 Local Context Size
While the parallel decoding mechanism theoretically allows the model to attend to all tokens, we aim to investigate whether the principle that a larger local context size equates to better performance holds true for FocusLLM as well. The results are shown in the last row of Table 5. As we reduce the local context size from 3.5K to 1K, it can be observed that performance for the majority of tasks experiences a slight decline, with the exception of the passkey retrieval task, which has very clear and concise instructions. This indicates that a larger local context size indeed provides higher quality local information.
Figure 7: Perplexity under different chunk size with the total sequence length fixed as 8K on three datasets.
4.5 Ablation Studies
We employ both Continuation Loss and Repetition Loss for the training of FocusLLM. The motivation behind this is to equip the model with the natural language modeling capability while also enhancing its ability to recover global information. Ablation studies on both losses, as detailed in Table 5, reveal that relying solely on the Continuation Loss enables the model to manage some tasks effectively. Nonetheless, for tasks with substantial dependencies on the preceding context, like HotpotQA and Retrieve.PassKey, the model’s efficacy deteriorates. Similarly, while employing the Repetition Loss ensures accurate restatement of the preceding context, the lack of generalizability of generating new tokens leads to a considerable decrease in performance. Therefore, the combined use of both loss functions is crucial for enhancing the performance and generalizability of FocusLLM.
5 Related Work
5.1 Long-context language models
Recent advancements in long-context modeling have seen a surge in innovative approaches that aim to transcend the limitations of transformer architectures. One research direction involves length extrapolation in transformers Peng et al. (2023); Jin et al. (2024), where methods like positional interpolation help models adapt to longer sequences Chen et al. (2023a). However, these techniques often fail to address the distraction issue caused by noisy content within extended texts Tworkowski et al. (2024). Another research branch focus on modifying the attention mechanism or employing compression techniques to maintain long texts within manageable lengths Chevalier et al. (2023); Zhang et al. (2024a). For instance, Xiao et al. (2023) discovered that retaining ‘sink tokens’ in conjunction with a sliding window can achieve smooth streaming output. Zhang et al. (2024a) expanded the context dramatically through compression. However, these methods share a common limitation: they cannot utilize information from all tokens.
5.2 Memory-enhanced Model
The integration of memory layers within transformer architectures has become a pivotal strategy for enhancing long-context comprehension Bertsch et al. (2024); Tworkowski et al. (2024); Fang et al. (2024). Common methodologies in memory-enhanced models often employ recurrent strategies that iteratively integrate information from the current window into a persistent memory Munkhdalai et al. (2024). Another approach is to initially encode the complete long text into memory tokens, which is then queried in to retrieve pertinent information as needed Xiao et al. (2024). For example, Yen et al. (2024) employ a small encoder to sequentially encode long text segments, followed by the integration of these encoded chunks into a decoder. However, the drawback of such methods is that the memory length does not extrapolate well, and expanding the memory still incurs substantial computational costs. In contrast, FocusLLM offers superior training efficiency and remains effective on exceedingly long texts.
6 Conclusion
In this work, we introduced FocusLLM, a novel framework that significantly extends the context length of LLMs. The core innovation lies in the parallel decoding strategy, which distribute the burden of understanding long texts across each chunk and effectively aggregating global information. FocusLLM stands out due to its remarkable training efficiency, allowing us to achieve substantial gains in context comprehension with minimal computational and memory cost. Compared to existing methods, FocusLLM not only exhibits superior performance across downstream tasks but also maintains low perplexities when handling extensive texts, up to 400K tokens. We hope FocusLLM can be an inspiring work for the community, driving further exploration of long-context models.
Appendix A More discussion
A.1 Architecture
In this section, we provide an in-depth explanation of the parallel decoding utilized by FocusLLM. As we mentioned in section 2.1.3, the processing of each chunk in FocusLLM can be considered a form of implicit decoding. This is because our goal is not to obtain representations for entire chunks of text, but rather to preserve information about the next token (based on the current chunk) through a special candidate token.
Furthermore, we can observe that the overall structure of FocusLLM shares similarities with the recently proposed concept of the decoder-decoder model architecture Sun et al. (2024), which differs from the traditional encoder-decoder model. Specifically, encoder-decoder models, like T5 Raffel et al. (2020), utilize a bidirectional encoder for input encoding and a unidirectional decoder for output generation. These models necessitate the re-encoding of the entire set of input and output tokens for each subsequent generation step, making them less suitable for autoregressive generation. In contrast, the decoder-decoder model (also FocusLLM), by caching the previously computed key/value vectors, enables the model to reuse them in the current generation step.
A.2 Loss Design
In the loss design of FocusLLM, in addition to using the commonly employed Continuation Loss, we designed Repetition Loss to explicitly train the model to maintain attention to the context. We believe that Repetition Loss is the reason why FocusLLM can achieve good results with a relatively small number of training tokens. Previous work has demonstrated that synthetic data can effectively enhance model capabilities or reduce costs Wu et al. (2024); Li et al. (2022, 2023b, 2024b). This has inspired us to design different types of synthetic data in the future to help FocusLLM quickly learn specific abilities.
Appendix B Details of Benchmarks
B.1 LongBench
LongBenchBai et al. (2023) includes 14 English tasks, 5 Chinese tasks, and 2 code tasks, with the average length of most tasks ranging from 5K to 15K. In experiments, we only utilize the English tasks. Detailed statistics of the tasks used in our paper are shown in Table 6.
Table 6: Detailed statistics of the tasks used in our paper of LongBench.
Task Task Type Eval metric Avg len Language Sample
HotpotQA Multi-doc QA F1 9,151 EN 200
2WikiMultihopQA Multi-doc QA F1 4,887 EN 200
MuSiQue Multi-doc QA F1 11,214 EN 200
MultiFieldQA-en Single-doc QA F1 4,559 EN 150
NarrativeQA Single-doc QA F1 18,409 EN 200
Qasper Single-doc QA F1 3,619 EN 200
GovReport Summarization Rouge-L 8,734 EN 200
QMSum Summarization Rouge-L 10,614 EN 200
MultiNews Summarization Rouge-L 2,113 EN 200
TriviaQA Few shot F1 8,209 EN 200
SAMSum Few shot Rouge-L 6,258 EN 200
TREC Few shot Accuracy 5,177 EN 200
PassageRetrieval-en Synthetic Accuracy 9,289 EN 200
LCC Code Edit Sim 1,235 Python/C#/Java 500
RepoBench-P Code Edit Sim 4,206 Python/Java 500
Table 7: Detailed statistics of the tasks used in our paper of $\infty$-Bench.
Task Name Context Examples Avg Input Tokens Avg Output Tokens
En.MC Fake Book 229 184.4k 5.3
Code.Debug Code Document 394 114.7k 4.8
Code.Run Synthetic 400 75.2k 1.3
Math.Find Synthetic 350 87.9k 1.3
Retrieve.PassKey Synthetic 590 122.4k 2.0
Retrieve.Number Synthetic 590 122.4k 4.0
Retrieve.KV[2̂] Synthetic 500 89.9k 22.7
B.2 $\infty$-Bench
The benchmark Zhang et al. (2024b) comprises 12 unique tasks, each crafted to assess different aspects of language processing and comprehension in extended contexts. Detailed statistics of the tasks used in our paper are shown in Table 7.
Appendix C Details of the effective lengths of models in Table 3 and 4.
Not all models are capable of processing infinite text lengths. Therefore, we provide a clear explanation of the effective input length for each method in Table 3 and Table 4. Specifically: (i) For models with a finite context length, we truncate the inputs by only preserving the system prompts and the tail of inputs to simulate real-world applications with streaming inputs like Xiao et al. (2024). For instance, in Table 3, these models include Original (4K), LChat (32K), Vic-16K (16K), Yarn (128K), PI (128K), and NTK (128K). (ii) For other models, including StreamingLLM, InfLLM, LongLlama, CEPE, Activation Beacon, and our FocusLLM, the input can theoretically be of any length. So we input the entire sequence on the two benchmarks.